_  Nursing_majors  Non-nursing_majors  Total
Males  96 1013  1109
Females 700  1724  2424
Total  796  2737  3533
(a) Find the probability that the student is male or a nursing major.
​P(being male or being nursing ​major)= __
​(Round to the nearest thousandth as​ needed.)

(b) Find the probability that the student is female or not a nursing major.
​P(being female or not being a nursing ​major)=__
​(Round to the nearest thousandth as​ needed.)
​(c) Find the probability that the student is not female or a nursing major.
​P(not being female or being a nursing ​major)=__
​(Round to the nearest thousandth as​ needed.)
​(d) Are the events​ "being male" and​ "being a nursing​ major" mutually​ exclusive? Explain.
A.
​No, because one​ can't be male and a nursing major at the same time.
B.
​Yes, because one​ can't be male and a nursing major at the same time.
C.
​No, because there are 96 males majoring in nursing.
D.
​Yes, because there are 96 males majoring in nursing.

Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

_ Nursing_majors Non-nursing_majors Total
Males 96 1013 1109
Females 700 1724 2424
Total 796 2737 3533
Male or a nursing major is 1109+796-both (96)=1809/0.512
Female or not a nursing major is 2424+2737-1724 (double counting)=3437/3533 (only male nursing majors are omitted. 0.973
Not a female or a nursing major leaves male non-nursing majors only, which is 1013/3533=0.287
Male and being a nursing major are not exclusive, because one can be a male and a nursing major. So the answer is C, No, they aren't mutually exclusive, because there are 96 males majoring in nursing.

Answer by MathTherapy(10552)

_     Nursing_majors           Non-nursing_majors       Total
Males   96                          1013                1109
Females 700                         1724                2424
Total   796                         2737                3533    

(a) Find the probability that the student is male or a nursing major.
​P(being male or being nursing ​major)= __
​(Round to the nearest thousandth as​ needed.)
​
(b) Find the probability that the student is female or not a nursing major.
​P(being female or not being a nursing ​major)=__
​(Round to the nearest thousandth as​ needed.)

​(c) Find the probability that the student is not female or a nursing major.
​P(not being female or being a nursing ​major)=__
​(Round to the nearest thousandth as​ needed.)

​(d) Are the events​ "being male" and​ "being a nursing​ major" mutually​ exclusive? Explain.

A.
​No, because one​ can't be male and a nursing major at the same time.

B.
​Yes, because one​ can't be male and a nursing major at the same time.

C.
​No, because there are 96 males majoring in nursing.

D.
​Yes, because there are 96 males majoring in nursing.

Notice that c) is the same as a).
Is this question from the Chamberlain College of Nursing Statistics course?